# Learning 3D Scene Analogies with Neural Contextual Scene Maps

### v1

@@ Abstract
Understanding scene contexts is crucial for machines to perform tasks and adapt prior knowledge in unseen or noisy 3D environments. As data-driven learning is intractable to comprehensively encapsulate diverse ranges of layouts and open spaces, we propose teaching machines to identify relational commonalities in 3D spaces. Instead of focusing on point-wise or object-wise representations, we introduce 3D scene analogies, which are smooth maps between 3D scene regions that align spatial relationships. Unlike well-studied single instance-level maps, these scene-level maps smoothly link large scene regions, potentially enabling unique applications in trajectory transfer in AR/VR, long demonstration transfer for imitation learning, and context-aware object rearrangement. To find 3D scene analogies, we propose neural contextual scene maps, which extract descriptor fields summarizing semantic and geometric contexts, and holistically align them in a coarse-to-fine manner for map estimation. This approach reduces reliance on individual feature points, making it robust to input noise or shape variations. Experiments demonstrate the effectiveness of our approach in identifying scene analogies and transferring trajectories or object placements in diverse indoor scenes, indicating its potential for robotics and AR/VR applications.
@@ 3.2.1 Training Descriptor Fields
As shown in Figure 5, we propose an automated procedure to generate positive and negative scene pairs. Our pipeline assumes a source dataset consisting of 3D scenes $\mathcal{D}_{\text{src}}{=}\{S_{i}\}$ with known object poses. Among the many possible definitions for a “correct" correspondence (e.g., appearance [94], style [67], or semantics [71]), we target finding point matches that share common nearby object semantics and local geometry, inspired from works in semantic correspondence [71, 72]. Based on this notion, the positive pairs $(S_{i},S^{+}_{i})$ are generated by swapping objects in each scene $S_{i}$ with randomly selected objects sharing the same semantic label from other scenes $\mathcal{D}_{\text{src}}\setminus S_{i}$. Here, the objects for replacement are sampled from the top-K (=100) list of objects having the most similar aspect ratios. Next, the negative pairs $(S_{i},S^{-}_{i})$ are generated by adding noise perturbations to the object poses, similar to LEGO-Net [117]. Note that we constrain the pose noise to planar translation and z-axis rotation to prevent floating objects or ground penetrations. The resulting triplet dataset $\mathcal{D}_{\text{triplet}}{=}\{(S_{i},S^{+}_{i},S^{-}_{i})\}$ is used for training.
@@ 4 Experiments
Metric PCP Bijectivity PCP Chamfer Acc.
@@ 5 Conclusion
We acknowledge several limitations that invite future work. Currently, our method outputs a single mapping, whereas generating multiple plausible mappings could better account for symmetries and multi-modal correspondences. Additionally, while our method handles a wide range of spatial variations, it may struggle when object positions swap, as such changes disrupt the initial affine mapping. Future work could explore more flexible alignment strategies to address these cases. Finally, our evaluation is based on semantic and local geometric similarity, but different tasks may require alternative notions of “correctness." Expanding the evaluation framework to incorporate task-specific criteria could provide deeper insights. Further details are in the supplementary material.
@@ B.6 Performance Analysis with Respect to the Number of RoI Points
Metric (Points Sampled per Object) PCK (50) PCK (100) PCK (200) PCK (400)
@@ C.3 Scene Pair Preparation for Evaluation
Recall these scene pairs contain pseudo ground-truth annotations for evaluating point-level accuracy of scene analogy estimations. For each scene, we first randomly select an object and its k-nearest neighbors (where k is randomly sampled from $\{2,\dots,4\}$). The points sampled from the selected objects are used as the region of interest $P_{\text{RoI}}$. Then, for objects not selected, we either randomly remove them by a probability of 0.5 or apply pose perturbation. Here translation noise is sampled from the uniform distribution $\mathcal{U}(-0.05,0.05)$ and rotation noise is obtained from the set of z-axis rotations with rotation angles sampled from $\mathcal{U}(-10^{\circ},10^{\circ})$. Next, we randomly add $N_{\text{add}}$ objects to open spaces in the scene (where $N_{\text{add}}$ is sampled from $\mathcal{U}(2,5)$). In this step, we retrieve the scene in the evaluation dataset with the closest object semantic label histogram, and select objects from that scene for addition. The objects are added by computing an occupancy grid map of the current scene and randomly choosing from collision-free locations [92]. Finally, we replace each object that has not been added or removed during the previous steps with a randomly selected object of the same semantic class, similar to training data generation explained in Section A.1. The resulting procedurally generated scenes contain realistic object placements while preserving meaningful object group structures for evaluation.
@@ C.4 Evaluation Metric Details
where $\alpha$ is a threshold parameter. For scene pairs labeled as unmatchable, the Chamfer accuracy is set to $1$ if no maps are produced, and 00 otherwise.
@@ Appendix D Limitations and Future Work
Handling Various Notions of “Correct" Correspondences
Inspired from prior works in semantic correspondence [71, 72, 44, 129], our work considers points having similar nearby object semantics and local geometry to be correct matches, and the descriptor fields are trained to support this notion of “correctness". We have demonstrated in Section 4.2 that this definition is useful for tasks such as trajectory transfer in robotics or object placement transfer in AR/VR. Nevertheless, we acknowledge that multiple definitions of "correct" correspondences exist depending on the task. For example, one may want to find scene analogies based on other attributes such as affordance (e.g., mapping ‘sittable’ areas from one scene to another) or appearance (e.g., matching furniture groups with a specific style). Due to the modular design of our approach of separating descriptor extraction and map estimation based on classical optimization, our method can be flexibly modified to handle such definitions of correctness. Specifically, one may train new descriptor sets for different correctness definitions and subsequently apply the map estimation process that does not require training.

### v2

@@ Abstract
Understanding scene contexts is crucial for machines to perform tasks and adapt prior knowledge in unseen or noisy 3D environments. As data-driven learning is intractable to comprehensively encapsulate diverse ranges of layouts and open spaces, we propose teaching machines to identify relational commonalities in 3D spaces. Instead of focusing on point-wise or object-wise representations, we introduce 3D scene analogies, which are smooth maps between 3D scene regions that align spatial relationships. Unlike well-studied single instance-level maps, these scene-level maps smoothly link large scene regions, potentially enabling unique applications in trajectory transfer in AR/VR, long demonstration transfer for imitation learning, and context-aware object rearrangement. To find 3D scene analogies, we propose neural contextual scene maps, which extract descriptor fields summarizing semantic and geometric contexts, and holistically align them in a coarse-to-fine manner for map estimation. This approach reduces reliance on individual feature points, making it robust to input noise or shape variations. Experiments demonstrate the effectiveness of our approach in identifying scene analogies and transferring trajectories or object placements in diverse indoor scenes, indicating its potential for robotics and AR/VR applications. Project page including the code is available through this link: https://82magnolia.github.io/3d_scene_analogies/.
@@ 3.2.1 Training Descriptor Fields
As shown in Figure 5, we propose an automated procedure to generate positive and negative scene pairs. Our pipeline assumes a source dataset consisting of 3D scenes $\mathcal{D}_{\text{src}}{=}\{S_{i}\}$ with known object poses. Among the many possible definitions for a “correct” correspondence (e.g., appearance [94], style [67], or semantics [71]), we target finding point matches that share common nearby object semantics and local geometry, inspired from works in semantic correspondence [71, 72]. Based on this notion, the positive pairs $(S_{i},S^{+}_{i})$ are generated by swapping objects in each scene $S_{i}$ with randomly selected objects sharing the same semantic label from other scenes $\mathcal{D}_{\text{src}}\setminus S_{i}$. Here, the objects for replacement are sampled from the top-K (=100) list of objects having the most similar aspect ratios. Next, the negative pairs $(S_{i},S^{-}_{i})$ are generated by adding noise perturbations to the object poses, similar to LEGO-Net [117]. Note that we constrain the pose noise to planar translation and z-axis rotation to prevent floating objects or ground penetrations. The resulting triplet dataset $\mathcal{D}_{\text{triplet}}{=}\{(S_{i},S^{+}_{i},S^{-}_{i})\}$ is used for training.
@@ 4 Experiments
Metric
PCP
Bijectivity PCP
Chamfer Acc.
@@ 5 Conclusion
We acknowledge several limitations that invite future work. Currently, our method outputs a single mapping, whereas generating multiple plausible mappings could better account for symmetries and multi-modal correspondences. Additionally, while our method handles a wide range of spatial variations, it may struggle when object positions swap, as such changes disrupt the initial affine mapping. Future work could explore more flexible alignment strategies to address these cases. Finally, our evaluation is based on semantic and local geometric similarity, but different tasks may require alternative notions of “correctness.” Expanding the evaluation framework to incorporate task-specific criteria could provide deeper insights. Further details are in the supplementary material.
Acknowledgements
This work was supported by the National Research Foundation of Korea(NRF) grant (No. RS-2023-00218601) and Institute of Information & communications Technology Planning & Evaluation (IITP) grant (No.RS-2021-II212068, Artificial Intelligence Innovation Hub) funded by the Korea government(MSIT), and the BK21 FOUR program of the Education and Research Program for Future ICT Pioneers, Seoul National University in 2025.
@@ B.6 Performance Analysis with Respect to the Number of RoI Points
Metric
(Points Sampled per Object)
PCK (50) PCK (100) PCK (200) PCK (400)
@@ C.3 Scene Pair Preparation for Evaluation
[…]
@@ C.4 Evaluation Metric Details
where $\alpha$ is a threshold parameter. For scene pairs labeled as unmatchable, the Chamfer accuracy is set to 111 if no maps are produced, and 0 otherwise.
@@ Appendix D Limitations and Future Work
Handling Various Notions of “Correct” Correspondences
Inspired from prior works in semantic correspondence [71, 72, 44, 129], our work considers points having similar nearby object semantics and local geometry to be correct matches, and the descriptor fields are trained to support this notion of “correctness”. We have demonstrated in Section 4.2 that this definition is useful for tasks such as trajectory transfer in robotics or object placement transfer in AR/VR. Nevertheless, we acknowledge that multiple definitions of ”correct” correspondences exist depending on the task. For example, one may want to find scene analogies based on other attributes such as affordance (e.g., mapping ‘sittable’ areas from one scene to another) or appearance (e.g., matching furniture groups with a specific style). Due to the modular design of our approach of separating descriptor extraction and map estimation based on classical optimization, our method can be flexibly modified to handle such definitions of correctness. Specifically, one may train new descriptor sets for different correctness definitions and subsequently apply the map estimation process that does not require training.